OpenStudy (aye_my_peeps):

mia borrowed $12000 at 5% interest for 6 years. what was the total interest mia paid???

OpenStudy (vijeya3):

hey mimi:) Do you know the formula for SI?

OpenStudy (aye_my_peeps):

ayee and not really......

OpenStudy (vijeya3):

Okay...doesn't matter. I'll tell ya SI is the Simple Interest.There are basically two types-Simple and Compound Interest.You can google them if you wanna know them in detail or even ask me about it:) Right now,let us concentrate on SI only.Okay?

OpenStudy (aye_my_peeps):

k

OpenStudy (vijeya3):

The formula for finding the SI is SI=R*P*T/100 where, P=Principal that is the amount at hand R=Rate at which we are gonna be paying it per year T=Time which here is the number of years. Can you do it now?:)

OpenStudy (vijeya3):

@AYE_MY_PEEPS ?

OpenStudy (aye_my_peeps):

no...

OpenStudy (vijeya3):

Okay so u didn't understand it?or you want me to do it for ya?

OpenStudy (aye_my_peeps):

both hahaha but mostly bc i didnt understand @Vijeya3

OpenStudy (bournville):

maybe i could do it for ya?

OpenStudy (bournville):

she seems off

OpenStudy (vijeya3):

Well..sorry was afk Okay,just remember the simple interest formula,alright? SI=P*R*T/100 Here, P=12000 R=5 T=6 Try now @AYE_MY_PEEPS ?

OpenStudy (aye_my_peeps):

uhhh

OpenStudy (vijeya3):

so what's ur answer?

OpenStudy (aye_my_peeps):

...

OpenStudy (vijeya3):

want me to do it?

OpenStudy (vijeya3):

which part do you not get?tell me?